OpenStudy (anonymous):

Function form help? 1. What's y-5=-2(x+1)^2 in intercept form? 2. What's y= (x+2)(x-3) in vertex and standard form?

OpenStudy (anonymous):

1.y=-2x^2-4x+3 2. y=x^2-x-6 (standard), y=(x-0.5)^2-25/4 (vertex)